how to solve √sē-20s+100? I know the answer is |s-10| but I need the steps in solving this. Somehow the -20s dissapears. Where does it go and why?
Answer by jim_thompson5910(35256)

s^2 - 20s + 100

s^2 - 10s - 10s + 100

s(s-10) - 10(s-10)

(s-10)(s-10)

(s-10)^2

So s^2 - 20s + 100 = (s-10)^2

This means

sqrt( s^2 - 20s + 100 ) = sqrt( (s-10)^2 )

sqrt( s^2 - 20s + 100 ) = |s-10|